OpenStudy (hybrik):

What is the inverse of y=2x+7, is it y=(x+7)/2?

OpenStudy (hybrik):

@ganeshie8 @welshfella

Nnesha (nnesha):

\[\huge\rm \color{reD}{y}=2\color{blue}{x}+7\] switch x and y to find inverse \[\huge\rm \color{blue}{x}=2\color{red}{y}+7\] now solve for y

OpenStudy (hybrik):

yeah, which is (x+y)/2

Nnesha (nnesha):

no solve for `y` what would be ur first step ?

OpenStudy (hybrik):

oops, x+7, typo

Nnesha (nnesha):

still.... wrong.

Nnesha (nnesha):

what was your first step ?

Nnesha (nnesha):

ello ??????...

OpenStudy (hybrik):

huh

OpenStudy (hybrik):

my answer was (x+7)/2 = y

Nnesha (nnesha):

that's wrong.

Nnesha (nnesha):

how did you get that ? what was ur first step ? show ur work maybe i'll able to find out the mistake...

OpenStudy (hybrik):

y-7=2x

Nnesha (nnesha):

yes right divide both sides by 2 and then switch x and y

OpenStudy (hybrik):

(y-7)/2 = x, (x-7)/2 = y

Nnesha (nnesha):

looks good!